OpenStudy (chacha1234):

How many solutions does this system have? y=3x-5 y=-x+4

rebeccaxhawaii (rebeccaxhawaii):

x^1 so only one

rebeccaxhawaii (rebeccaxhawaii):

if it were x^2 then it would have 2

OpenStudy (chacha1234):

y=-3x -9x-4y=21

RhondaSommer (rhondasommer):

well you can replace y with what it equals.\[y=3x-5\]\[y=-x+4\]\[-x+4=3x+5\]\[4=2x+5\]\[-1=2x\]\[x=\frac{-1}{2}\]

RhondaSommer (rhondasommer):

then, you have to place \[-0.5~or~\frac{-1}{2}\] into \[y=3x-5\]in place of x. so \[y=3(-0.5)-5\]can you simplify that?